OpenStudy (anonymous):

2 ferryboats started at the same instant from opposite sides of a river, traveling across the water at right angles to the shore. They first met 720 yards from the nearest shore and, on reaching the opposite side, spent 10 minutes in the slip before starting back. They met again 400 yards from the nearest shore. What is the width of the river?

OpenStudy (anonymous):

If the first boat has speed v1 and the second has v2. Let the width be x. The first boat has travelled 720m. The second x-720. They have travelled for same time so 720/v1=(x-720)/v2. v2/v1=(x-720)/720. Similarly the second time, 400/v2=(x-400)/v1. v2/v1=400/x-400. so (x-720)(x-400)=720*400. Solve the quadratic.

OpenStudy (anonymous):

could you elaborate it further? it's complicated. . .

OpenStudy (anonymous):

pleaseee.

OpenStudy (y2o2):

If there velocity were the same in the i the two times , they would have met at the same distance from the nearest shore in the two times.

OpenStudy (anonymous):

Ok.  So here velocity = distance/time. The time they travel to meet each other is same as they started out at the same time.